OpenStudy (anonymous):

i am in 4th grade i need a refresh, someone to show me and explain the problem so that i can solve other problems...2/8+1/4+?=1 can you please show you got the answer please.. thxs

OpenStudy (anonymous):

2/8 = 1/4 1/4 + 1/4 = 1/2 1/2 + ? = 1 ? = 1/2

OpenStudy (anonymous):

ok so i see now how you got the answer, now what if it's an odd number like 2/5+5/10+?= did i do this problem right so its 1/4+3/4+?=1 i put 0 is that right?

OpenStudy (anonymous):

1/10 is answer right hope so cu thats final answer for me